OpenStudy (anonymous):

write an equation of the line with slope 3/7 containing the point (3,8)

OpenStudy (anonymous):

y = mx + b since slope is given that will be the value for m so now you have y = (3/7)x + b now you plug in your values for x and y 8 = (3/7)3 + b b = 8 - 9/7 = -17/10 so your equation will be y = (3/7)x - 17/10

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

no problems